# Computation of attractor dimension and maximal sums of Lyapunov exponents using polynomial optimization

### v1

@@ 2.1 Tangent vectors, Lyapunov exponents and Lyapunov dimension
Since $\mu_{j+1}=M_{j}-M_{j+1}>0$, rearranging the constraint in 22 gives the constraint in 19 as an equivalent form. This infimum is in $[0,1)$ and so is unchanged by the $B\in[0,1)$ constraint in 19. Thus expression 19 for $d_{L}$ is equivalent to expression 18 in the definition. ∎
@@
for all $x_{0}\in\mathcal{B}$, then the Lyapunov dimension is bounded above by $d_{L}\leq k+B$. Note that 23 cannot hold for all $x_{0}$ unless $B\geq 0$.
@@ 2.2 Sphere projection approach
Our first approach to bounding the LE sums $M_{k}$ of 2.2 uses the formulation 5 for bounding infinite-time averages. This is a generalization of the approach in [34], which was described only for the $k=1$ case. First note that the time average in the definition of $M_{k}$ can be expressed as a time integral along trajectories in the augmented state space for $x(t)$ and $y^{k}(t)$:
@@
where $(y^{k})^{\mathsf{T}}$ is the transpose of the vector $y^{k}$. The formulation 5 for bounding time averages can be applied to the quantity in 26 only if each $(x,y^{k})$ remains bounded forward in time, but $y^{k}(t)$ will be unbounded when $M_{k}$ is positive. The same tangent space dynamics can be captured with bounded variables by projecting $y^{k}$ onto a sphere, by introducing $z^{k}=y^{k}/|y^{k}|$. Expressing the right-hand side of 26 in terms of $z^{k}$ and then maximizing both sides over the initial tangent space directions $z^{k}_{0}$ gives
$M_{k}(x_{0})=\sup_{z_{0}^{k}\in\mathbb{S}^{n_{k}-1}}\overline{\Phi_{k}},$ (27)
@@
along trajectories in the $(x,z^{k})$ state space. The ODE governing $z^{k}(t)$ follows from the ODE 15 for $y^{k}(t)$. The $x(t)$ dynamics together with the projected dynamics of $k$-vectors in its tangent space gives the ODE system
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}x\\ z^{k}\end{bmatrix}=\begin{bmatrix}f(x)\\ \ell_{k}(x,z^{k})\end{bmatrix}$, (29)
on the augmented state space $\mathcal{B}\times\mathbb{S}^{n_{k}-1}$, where
@@
Representations of $M_{k}$ as time averages over the projected tangent space dynamics 29 have often appeared in the study of stochastic dynamics, where they are called Furstenberg–Khasminskii formulas [2, chapter 6]. (In the stochastic case, $\Phi_{k}$ is defined with additional terms and is averaged with respect to stationary measures.)
Auxiliary function methods for bounding time averages can be applied to $M_{k}$ by using its representation 27 as a time average in the augmented state space for $(x,z^{k})$. In particular, we directly apply the formula 5 with state variable $X=(x,z^{k})$ on domain $\mathcal{B}\times\mathbb{S}^{n_{k}-1}$ and with $F(X)$ being the right-hand side of 29. The resulting inequality is the first part of proposition 2.6 below, for which no further proof is needed. If $F\in C^{1}$ and the domain is compact, then 5 holds with equality, as proved in [46]. This yields the second part of proposition 2.6 since assuming $f\in C^{2}$ gives that $F\in C^{1}$. The $k=1$ special case of proposition 2.6 is proposition 1 in [34].
@@
Let $\frac{\mathrm{d}}{\mathrm{d}t}x(t)=f(x(t))$ with $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$ and dynamics forward invariant on $\mathcal{B}\subset\mathbb{R}^{n}$. Let $\Phi_{k}(x,z^{k})$ and $\ell_{k}(x,z^{k})$ be defined as in 28 and 30, respectively.
@@
    $\sup_{x_{0}\in\mathcal{B}}M_{k}(x_{0})\leq\inf_{V\in C^{1}\left(\mathcal{B}\times\mathbb{R}^{n_{k}}\right)}\sup_{\begin{subarray}{c}x\in\mathcal{B}\hskip 13.0pt\\ z^{k}\in\mathbb{S}^{{n_{k}}-1}\end{subarray}}\left[\Phi_{k}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{k}}V\right],$ (31)
@@
The Lyapunov dimension 18 can be bounded above using similar ideas. One way is to apply upper bounds $M_{k}\leq B_{k}$ that have already been found using the right-hand side of 31 or its SOS relaxations. If such bounds are computed with $B_{k}\geq 0$ but $B_{k+1}<0$ for some $k\in\{1,\ldots,n\}$, then proposition 2.7 below gives an upper bounds on $d_{L}$. To apply this result, one does not need to know whether $k$ is the minimal integer $j$ appearing in the definition 18 of $d_{L}$. It is enough to know that $B_{k+1}<0$, which implies $k\geq j$.
@@
Suppose that, for some $k\in\{1,\ldots,n\}$ and values $B_{k}\geq 0$ and $B_{k+1}<0$,
@@
Then, the Lyapunov dimension 18 is bounded above by
@@
The meaning of $j$ in 2.3 and the fact that $B_{k+1}<0$ together imply $k\geq j$. According to 2.5, it suffices for the inequality 23 to hold for all $x_{0}$ with $B=\frac{B_{k}}{B_{k}-B_{k+1}}$. For any $x_{0}\in\mathcal{B}$, the assumptions give $M_{k}(x_{0})\leq B_{k}$ and $0<-B_{k+1}\leq-M_{k+1}(x_{0})$, so
@@
which is the required inequality 23. This concludes the proof. ∎
The upper bound on Lyapunov dimension given by proposition 2.7 cannot be sharp for certain systems, even if $B_{k}$ and $B_{k+1}$ are each sharp bounds in 32. The reason is that the inequality 34, when maximized over $x_{0}$, is an equality only if $M_{k}(x_{0})$ and $M_{k+1}(x_{0})$ attain their maxima on the same orbits. This is not true in general; see section 3.2 for a counterexample. We thus give a second formulation for upper bounds on $d_{L}$ that is generally sharper than proposition 2.7. This formulation directly considers the ratio in the Kaplan–Yorke formula 18, rather than separately bounding the $M_{k}$ and $M_{k+1}$ terms in that ratio. In particular, we consider the $x(t)$ dynamics together with the projected dynamics of $k$-vectors and $(k+1)$-vectors in its tangent space,
@@
where $\ell_{k}$ is defined by 30, and likewise for $\ell_{k+1}$. The augmented state space of this system is $(x,z^{k},z^{k+1})\in\mathcal{B}\times{S}^{n_{k}-1}\times\mathbb{S}^{n_{k+1}-1}$. The constraint in 19 from 2.4 can then be expressed in terms of time averages in the ODE system 37. For this we define
@@
$\overline{\Psi}^{B}(x_{0},z^{k}_{0},z^{k+1}_{0})\leq 0~~\text{for all}~x_{0}\in\mathcal{B},~z^{k}_{0}\in\mathbb{S}^{n_{k}-1},~z^{k+1}_{0}\in\mathbb{S}^{n_{k+1}-1},$ (39)
where the time average denoted by the overbar is along trajectories of 37. Condition 39 is equivalent to the first condition in 19 of 2.4, as explained below in the proof of proposition 2.8. The condition 39 can be enforced using the general formulation 5 for bounding time averages, applied to the augmented system 37. This yields the first part of proposition 2.8, whose sharpness is addressed by its second part. Proposition 2.8 can be applied only when the LE sum $M_{k+1}$ for certain $k$ is known to be negative on every trajectory. Such negativity can be confirmed by using 31, or its SOS relaxation that we introduce in section 2.4, to obtain a negative upper bound on $M_{k+1}$.
@@
Let $\frac{\mathrm{d}}{\mathrm{d}t}x(t)=f(x(t))$ with $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$ and dynamics forward invariant on $\mathcal{B}\subset\mathbb{R}^{n}$. Let $\Phi_{k}(x,z^{k})$ and $\ell_{k}(x,z^{k})$ be as defined by 28 and 30, and likewise for $\Phi_{k+1}(x,z^{k+1})$ and $\ell_{k+1}(x,z^{k+1})$.
@@
    Suppose each trajectory $x(t)$ in $\mathcal{B}$ is bounded for $t\geq 0$. Let the positive integer $k\leq n$ be such that $\sup_{x_{0}\in\mathcal{B}}M_{k+1}(x_{0})<0$. Then the global Lyapunov dimension $d_{L}$ over $\mathcal{B}$ is bounded above by
    $d_{L}\leq\inf\limits_{\begin{subarray}{c}B\in[0,1]\\ V\in C^{1}~~\end{subarray}}(k+B)\quad\text{s.t.}\quad\begin{array}[t]{l}\Psi^{B}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{k}}V+\ell_{k+1}\cdot D_{z^{k+1}}V\leq 0\\[4.0pt] \text{for all}\quad x\in\mathcal{B},~z^{k}\in\mathbb{S}^{n_{k}-1},~z^{k+1}\in\mathbb{S}^{n_{k+1}-1},\end{array}$ (40)
    where $V:\mathcal{B}\times\mathbb{S}^{{n_{k}}-1}\times\mathbb{S}^{n_{k+1}-1}\to\mathbb{R}$, and $D_{x}V$, $D_{z^{k}}V$ and $D_{z^{k+1}}V$ denote the gradients of $V(x,z^{k},z^{k+1})$ with respect to each argument.
@@
    $B_{\varepsilon}=\inf\limits_{\begin{subarray}{c}B\in[0,1]\\ V\in C^{1}\hskip 7.0pt\end{subarray}}B\quad\text{s.t.}\quad\begin{array}[t]{l}\Psi^{B}+f\cdot D_{x}V+\ell_{j}\cdot D_{z^{j}}V+\ell_{j+1}\cdot D_{z^{j+1}}V\leq\varepsilon\\[4.0pt] \text{for all}\quad x\in\mathcal{B},~z^{j}\in\mathbb{S}^{n_{j}-1},~z^{j+1}\in\mathbb{S}^{n_{j+1}-1}.\end{array}$ (42)
@@
For the first part, the assumption that $\sup_{x_{0}\in\mathcal{B}}M_{k+1}(x_{0})<0$ means that $k\geq j$ where $j$ is as defined in 2.3. Applying 2.5, we therefore have
@@
$(1-B)\,\sup_{z_{0}^{k}\in\mathbb{S}^{n_{k}}}\overline{\Phi}_{k}(x_{0},z_{0}^{k})+B\,\sup_{z_{0}^{k+1}\in\mathbb{S}^{n_{k+1}}}\overline{\Phi}_{k+1}(x_{0},z_{0}^{k+1})~\leq~0\quad\forall x_{0}\in\mathcal{B}.$ (44)
This inequality holds if and only if it holds without the suprema for all $z_{0}^{k}$ and $z_{0}^{k+1}$, which is exactly the $\overline{\Psi}^{B}\leq 0$ condition 39. Writing this latter condition in place of the equivalent constraint in 43 gives
$d_{L}\leq\inf_{B\in[0,1)}(k+B)\quad\text{s.t.}\quad\overline{\Psi}^{B}(x_{0},z_{0}^{k},z_{0}^{k+1})\leq 0~~\text{for all}~(x_{0},z_{0}^{k},z_{0}^{k+1})\in\mathcal{B}\times\mathbb{S}^{n_{k}-1}\times\mathbb{S}^{n_{k+1}-1}.$ (45)
We want to replace the constraint in 45 with one that does not involve time averages. To do so we upper bound $\overline{\Psi}^{B}$ using the general formulation 5 for time averages by letting $X=(x,z^{k},z^{k+1})$ on domain $\Omega=\mathcal{B}\times\mathbb{S}^{n_{k}-1}\times\mathbb{S}^{n_{k+1}-1}$, and letting $F(X)$ be the right-hand side of 37. This gives
$\sup_{\begin{subarray}{c}x_{0}\in\mathcal{B}~~~~~\\ z^{k}_{0}\in\mathbb{S}^{n_{k}-1}\\ z^{k+1}_{0}\in\mathbb{S}^{n_{k+1}-1}\end{subarray}}\overline{\Psi}_{B}\leq\inf_{V\in C^{1}}\sup_{\begin{subarray}{c}x\in\mathcal{B}\\ z^{k}\in\mathbb{S}^{n_{k}-1}\\ z^{k+1}\in\mathbb{S}^{n_{k+1}-1}\end{subarray}}\left[\Psi^{B}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{k}}V+\ell_{k+1}\cdot D_{z^{k+1}}V\right].$ (46)
The constraint of 43, which is equivalent to nonpositivity of the left-hand side of 46, can be strengthened by requiring there to exist $V\in C^{1}$ for which the inner supremum on the right-hand side is nonpositive. This gives the first part of the proposition.
For the second part, note that 43 is an equality when $k=j$, according to 2.4. That equality can be expressed as
$d_{L}=\inf_{B\in[0,1)}(j+B)\quad\text{s.t.}\quad\overline{\Psi}^{B}(x_{0},z_{0}^{j},z_{0}^{j+1})\leq 0~~\text{for all}~(x_{0},z_{0}^{j},z_{0}^{j+1})\in\mathcal{B}\times\mathbb{S}^{n_{j}-1}\times\mathbb{S}^{n_{j+1}-1},$ (47)
where equivalence of the constraints in 43 and 47 is as in the first part. To prove the second inequality in the claim 41, let $B$ be any value for which the constraint in 47 holds. It suffices to prove existence of $V$ for which constraint in the definition 42 of $B_{\varepsilon}$ holds also. This will imply that the infimum over $B$ in 42 is at least as small as in 47.
Under the assumptions that $\mathcal{B}$ is compact and $f\in C^{2}$, the $X$ domain $\Omega$ is compact, and $F(X)$ defined as the right-hand side of 37 is $C^{1}$ on an open region containing $\Omega$. For such $X$ and $F$, the general upper bound 5 is an equality [46], so in particular 46 is an equality. Choosing $k=j$, equality in 47 implies that the constraint in 47 is equivalent to
$0\geq\inf_{V\in C^{1}}\sup_{\begin{subarray}{c}x\in\mathcal{B}\\ z^{k}\in\mathbb{S}^{n_{k}-1}\\ z^{k+1}\in\mathbb{S}^{n_{k+1}-1}\end{subarray}}\left[\Psi^{B}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{k}}V+\ell_{k+1}\cdot D_{z^{k+1}}V\right].$ (48)
@@
$\left[\Psi^{B}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{k}}V+\ell_{k+1}\cdot D_{z^{k+1}}V\right]\leq\varepsilon\quad\forall~(x,z^{k},z^{k+1})\in\mathcal{B}\times\mathbb{S}^{n_{k}-1}\times\mathbb{S}^{n_{k+1}-1}.$ (49)
(Such $V$ need not exist with $\varepsilon=0$ if the infimum over $V$ in the constraint of 46 is not attained.) Condition 49 is the same as the constraint in the minimization 42 defining $B_{\varepsilon}$, and this implies $B_{\varepsilon}\leq B$. We have shown that $B_{\varepsilon}\leq B$ for any $B$ satisfying the constraint in 47, and thus $d_{L}-j\leq B_{\varepsilon}$. This establishes the second inequality in 41.
@@
$M_{k}(x_{0})+B\mu_{k+1}(x_{0})\leq\varepsilon,$ (50)
which we divide by $\mu_{k+1}>0$ to find
@@
Since 52 holds at all $B$ where the constraint set in the definition 42 of $B_{\varepsilon}$ is not empty, minimizing over such $B$ gives 52 with $B_{\varepsilon}$ on the right-hand side. This establishes the first inequality in 41, which completes the proof. ∎

### v2

@@ 2.1 Tangent vectors, Lyapunov exponents and Lyapunov dimension
Since $\mu_{j+1}=M_{j}-M_{j+1}>0$, rearranging the constraint in 22 gives the constraint in 19 as an equivalent form. This infimum is in $[0,1)$ and so is unchanged by the $B\in[0,1)$ constraint in 19. Thus expression 19 for $d_{L}$ is equivalent to 18. ∎
@@
for all $x_{0}\in\mathcal{B}$, then $d_{L}\leq k+B$. Note that 23 cannot hold for all $x_{0}$ unless $B\geq 0$.
@@ 2.2 Sphere projection approach
Our first approach to bounding the LE sums $M_{k}$ uses the formulation 5 for bounding infinite-time averages. This is a generalization of the approach in [34], which was described only for the leading LE—i.e., the $k=1$ case. First note that the time average in the 2.2 for $M_{k}$ can be expressed as a time integral along trajectories in the augmented state space for $(x,y^{k})$:
@@
where $(y^{k})^{\mathsf{T}}$ is the transpose of the vector $y^{k}$. The formulation 5 for bounding time averages could be applied to the quantity in 26 if each $(x,y^{k})$ were to remain bounded for $t\geq 0$, but $y^{k}(t)$ will grow without bound when $M_{k}$ is positive. Instead, the tangent space dynamics can be captured with bounded variables by projecting $y^{k}$ onto a sphere, by introducing $z^{k}=y^{k}/|y^{k}|$. Expressing the right-hand side of 26 in terms of $z^{k}$ and then maximizing both sides over the initial tangent space directions $z^{k}_{0}$ gives
$M_{k}(x_{0})=\sup_{z_{0}^{k}\in\mathbb{S}^{C_{n}^{k}-1}}\overline{\Phi_{k}},$ (27)
@@
along trajectories in the $(x,z^{k})$ state space. The ODE for $z^{k}$ follows from the ODE 15 for $y^{k}$, and it couples with the ODE 6 for $x$ to form the augmented system
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}x\\ z^{k}\end{bmatrix}=\begin{bmatrix}f(x)\\ \ell_{k}(x,z^{k})\end{bmatrix}$ (29)
on state space $\mathcal{B}\times\mathbb{S}^{C_{n}^{k}-1}$, where
@@
Expressions for $M_{k}$ in terms of the projected tangent space dynamics 29 have often appeared in the study of stochastic dynamics, where they are called Furstenberg–Khasminskii formulas [1, chapter 6]. In the stochastic case, the formula 27 differs by additional terms in $\Phi_{k}$, and the averaging is with respect to stationary measures.
Auxiliary function methods for bounding time averages can be applied to $M_{k}$ by using its representation 27 as a time average in the augmented state space for $(x,z^{k})$. In particular, we directly apply the formula 5 with state variable $X=(x,z^{k})$ on domain $\mathcal{B}\times\mathbb{S}^{C_{n}^{k}-1}$ and with $F(X)$ being the right-hand side of 29. The resulting inequality is the first part of proposition 2.6 below, for which no further proof is needed. If $F\in C^{1}$ and the domain is compact, then 5 holds with equality, as proved in [50]. This yields the second part of proposition 2.6 since assuming $f\in C^{2}$ ensures $F\in C^{1}$. The $k=1$ special case of proposition 2.6 is proposition 1 in [34].
@@
Let $x(t)$ evolve by the ODE 6, where $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$ and $\mathcal{B}$ is forward invariant. Let $\Phi_{k}(x,z^{k})$ and $\ell_{k}(x,z^{k})$ be as defined by 28 and 30, respectively.
@@
    $\sup_{x_{0}\in\mathcal{B}}M_{k}(x_{0})\leq\inf_{V\in C^{1}\left(\mathcal{B}\times\mathbb{R}^{C_{n}^{k}}\right)}\sup_{\begin{subarray}{c}x\in\mathcal{B}\hskip 13.0pt\\ z^{k}\in\mathbb{S}^{{C_{n}^{k}}-1}\end{subarray}}\left[\Phi_{k}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{k}}V\right],$ (31)
@@
The Lyapunov dimension $d_{L}$ can be bounded above using similar ideas. One way is to apply upper bounds $M_{k}\leq B_{k}$ that have already been found using the right-hand side of 31 or its SOS relaxations. If such bounds are computed with $B_{k}\geq 0$ but $B_{k+1}<0$ for some integer $k$, then proposition 2.7 below gives an upper bound on $d_{L}$. To apply this result, one does not need to know whether $k$ is the minimal integer $j$ appearing in 2.3 for $d_{L}$. It is sufficient that $B_{k+1}<0$, which implies $k\geq j$.
@@
Suppose that, for some $k\in\{1,\ldots,n-1\}$ and values $B_{k}\geq 0$ and $B_{k+1}<0$,
@@
Then, the global Lyapunov dimension over $\mathcal{B}$ (2.3) is bounded above by
@@
The meaning of $j$ in 2.3 and the fact that $B_{k+1}<0$ together imply $k\geq j$. Due to 2.5, it suffices for the inequality 23 to hold for all $x_{0}$ with $B=\frac{B_{k}}{B_{k}-B_{k+1}}$. For any $x_{0}\in\mathcal{B}$, the assumptions give $M_{k}(x_{0})\leq B_{k}$ and $0<-B_{k+1}\leq-M_{k+1}(x_{0})$, so
@@
which is the required inequality 23. ∎
The upper bound on Lyapunov dimension given by proposition 2.7 cannot be sharp for certain systems, even if $B_{k}$ and $B_{k+1}$ are each sharp bounds in 32. The reason is that the inequality 34, when both sides are maximized over $x_{0}$, is an equality only if $M_{k}(x_{0})$ and $M_{k+1}(x_{0})$ attain their maxima on the same orbits. This is not true in general; see section 3.2 for a counterexample. We thus give a second formulation for upper bounds on $d_{L}$ that is generally sharper than proposition 2.7. This formulation directly considers the ratio in the Kaplan–Yorke formula 18, rather than separately bounding the $M_{k}$ and $M_{k+1}$ terms in that ratio. In particular, we consider the $x$ dynamics together with the projected dynamics of $k$-vectors and $(k+1)$-vectors,
@@
where $\ell_{k}$ and $\ell_{k+1}$ are defined by 30. The augmented state space of this system is $(x,z^{k},z^{k+1})\in\mathcal{B}\times\mathbb{S}^{C_{n}^{k}-1}\times\mathbb{S}^{C_{n}^{k+1}-1}$. The constraint in 19 from 2.4 can then be expressed in terms of trajectories of the ODE system 37. For this we define
@@
$\overline{\Psi}^{B}(x_{0},z^{k}_{0},z^{k+1}_{0})\leq 0~~\text{for all}~(x_{0},z^{k}_{0},z^{k+1}_{0})\in\mathcal{B}\times\mathbb{S}^{C_{n}^{k}-1}\times\mathbb{S}^{C_{n}^{k+1}-1},$ (39)
where the time average denoted by the overbar is along trajectories of 37. Condition 39 is equivalent to the condition in 19 of 2.4, as explained below in the proof of proposition 2.8. The condition 39 can be enforced using the general formulation 5 for bounding time averages, applied to the augmented system 37. This yields the first part of proposition 2.8, whose sharpness is addressed by its second part. Proposition 2.8 can be applied only when the LE sum $M_{k+1}$ for certain $k$ is known to be negative on every trajectory. Such negativity can be confirmed by using 31, or its SOS relaxation that we introduce in section 2.4, to obtain a negative upper bound on $M_{k+1}$.
@@
Let $x(t)$ evolve by the ODE 6, where $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$ and $\mathcal{B}$ is forward invariant. Let $\Phi_{k}(x,z^{k})$ and $\ell_{k}(x,z^{k})$ be as defined by 28 and 30, and likewise for $\Phi_{k+1}(x,z^{k+1})$ and $\ell_{k+1}(x,z^{k+1})$.
@@
    Suppose each trajectory $x(t)$ in $\mathcal{B}$ is bounded for $t\geq 0$. Let the positive integer $k<n$ be such that $\sup_{x_{0}\in\mathcal{B}}M_{k+1}(x_{0})<0$. Then the global Lyapunov dimension $d_{L}$ over $\mathcal{B}$ is bounded above by
    $d_{L}\leq\inf\limits_{\begin{subarray}{c}B\in[0,1)\\ V\in C^{1}~~\end{subarray}}(k+B)\quad\text{s.t.}\quad\begin{array}[t]{l}\Psi^{B}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{k}}V+\ell_{k+1}\cdot D_{z^{k+1}}V\leq 0\\[4.0pt] \text{for all}~~(x,z^{k},z^{k+1})\in\mathcal{B}\times\mathbb{S}^{C_{n}^{k}-1}\times\mathbb{S}^{C_{n}^{k+1}-1},\end{array}$ (40)
    where $V:\mathcal{B}\times\mathbb{S}^{{C_{n}^{k}}-1}\times\mathbb{S}^{C_{n}^{k+1}-1}\to\mathbb{R}$, and $D_{x}V$, $D_{z^{k}}V$ and $D_{z^{k+1}}V$ denote the gradients of $V(x,z^{k},z^{k+1})$ with respect to each argument.
@@
    $B_{\varepsilon}=\inf\limits_{\begin{subarray}{c}B\in[0,1)\\ V\in C^{1}\hskip 7.0pt\end{subarray}}B\quad\text{s.t.}\quad\begin{array}[t]{l}\Psi^{B}+f\cdot D_{x}V+\ell_{j}\cdot D_{z^{j}}V+\ell_{j+1}\cdot D_{z^{j+1}}V\leq\varepsilon\\[4.0pt] \text{for all}~~(x,z^{k},z^{k+1})\in\mathcal{B}\times\mathbb{S}^{C_{n}^{k}-1}\times\mathbb{S}^{C_{n}^{k+1}-1}.\end{array}$ (42)
@@
For the first part, the assumption that $\sup_{x_{0}\in\mathcal{B}}M_{k+1}(x_{0})<0$ means that $k\geq j$, where $j$ is as in 2.3. By 2.5, we therefore have
@@
$(1-B)\left[\sup_{z_{0}^{k}\in\mathbb{S}^{C_{n}^{k}}}\overline{\Phi}_{k}(x_{0},z_{0}^{k})\right]+B\left[\sup_{z_{0}^{k+1}\in\mathbb{S}^{C_{n}^{k+1}}}\overline{\Phi}_{k+1}(x_{0},z_{0}^{k+1})\right]~\leq~0~~\text{for all}~x_{0}\in\mathcal{B}.$ (44)
This condition is equivalent to the $\overline{\Psi}^{B}\leq 0$ condition 39, so the latter can replace the constraint in 43 to give
$d_{L}\leq\inf_{B\in[0,1)}(k+B)\quad\text{s.t.}\quad\overline{\Psi}^{B}(x_{0},z_{0}^{k},z_{0}^{k+1})\leq 0~~\text{for all}~(x_{0},z_{0}^{k},z_{0}^{k+1})\in\mathcal{B}\times\mathbb{S}^{C_{n}^{k}-1}\times\mathbb{S}^{C_{n}^{k+1}-1}.$ (45)
We want to replace the constraint in 45 with one that does not involve time averages. To do so we bound $\overline{\Psi}^{B}$ above using the general formulation 5 for time averages by letting $X=(x,z^{k},z^{k+1})$ on domain $\Omega=\mathcal{B}\times\mathbb{S}^{C_{n}^{k}-1}\times\mathbb{S}^{C_{n}^{k+1}-1}$, and letting $F(X)$ be the right-hand side of 37. This gives
$\sup_{\begin{subarray}{c}x_{0}\in\mathcal{B}\hskip 18.0pt\\ z^{k}_{0}\in\mathbb{S}^{C_{n}^{k}-1}\\ z^{k+1}_{0}\in\mathbb{S}^{C_{n}^{k+1}-1}\end{subarray}}\overline{\Psi}_{B}\leq\inf_{V\in C^{1}}\sup_{\begin{subarray}{c}x\in\mathcal{B}\hskip 14.0pt\\ z^{k}\in\mathbb{S}^{C_{n}^{k}-1}\\ z^{k+1}\in\mathbb{S}^{C_{n}^{k+1}-1}\end{subarray}}\left[\Psi^{B}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{k}}V+\ell_{k+1}\cdot D_{z^{k+1}}V\right].$ (46)
The constraint in 43, which is equivalent to nonpositivity of the left-hand side of 46, can be strengthened by requiring there to exist $V\in C^{1}$ for which the inner supremum on the right-hand side is nonpositive. This gives the first part of the proposition.
For the second part, note that 43 is an equality when $k=j$, according to 2.4, and likewise 45 is an equality:
$d_{L}=\inf_{B\in[0,1)}(j+B)\quad\text{s.t.}\quad\overline{\Psi}^{B}(x_{0},z_{0}^{j},z_{0}^{j+1})\leq 0~~\text{for all}~(x_{0},z_{0}^{j},z_{0}^{j+1})\in\mathcal{B}\times\mathbb{S}^{C_{n}^{j}-1}\times\mathbb{S}^{C_{n}^{j+1}-1}.$ (47)
To prove the second inequality in the claim 41, let $B$ be any value for which the constraint in 47 holds. It suffices to prove existence of $V$ for which constraint in the definition 42 of $B_{\varepsilon}$ holds also. This will imply that the infimum over $B$ in 42 is at least as small as in 47.
Under the assumptions that $\mathcal{B}$ is compact and $f\in C^{2}$, the $X$ domain $\Omega$ is compact, and $F(X)$, the right-hand side of 37, is $C^{1}$ on an open region containing $\Omega$. For such $X$ and $F$, the general upper bound 5 is an equality [50], so in particular 46 is an equality. Equality in 46 with $k=j$, and in 47, implies that the constraint in 47 is equivalent to
$0\geq\inf_{V\in C^{1}}\sup_{\begin{subarray}{c}x\in\mathcal{B}\\ z^{k}\in\mathbb{S}^{C_{n}^{k}-1}\\ z^{k+1}\in\mathbb{S}^{C_{n}^{k+1}-1}\end{subarray}}\left[\Psi^{B}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{k}}V+\ell_{k+1}\cdot D_{z^{k+1}}V\right].$ (48)
@@
$\left[\Psi^{B}+f\cdot D_{x}V+\ell_{k}\cdot D_{z^{k}}V+\ell_{k+1}\cdot D_{z^{k+1}}V\right]\leq\varepsilon\quad\forall~(x,z^{k},z^{k+1})\in\mathcal{B}\times\mathbb{S}^{C_{n}^{k}-1}\times\mathbb{S}^{C_{n}^{k+1}-1}.$ (49)
(Such $V$ need not exist with $\varepsilon=0$ if the infimum over $V$ in 46 is not attained.) Condition 49 is the same as the constraint in the minimization 42 defining $B_{\varepsilon}$, and this implies $B_{\varepsilon}\leq B$. We have shown that $B_{\varepsilon}\leq B$ for any $B$ satisfying the constraint in 47, and thus $j+B_{\varepsilon}\leq d_{L}$. This establishes the second inequality in 41.
@@
$(1-B)M_{k}(x_{0})+BM_{k+1}(x_{0})\leq\varepsilon.$ (50)
Dividing by the positive quantity $\mu_{k+1}=M_{k}-M_{k+1}$ gives
@@
This inequality holds at every $B$ for which there exists a $V$ satisfying the constraint in the minimization 42 that defines $B_{\varepsilon}$. Therefore, minimizing over such $B$ gives 52 with $B_{\varepsilon}$ on the right-hand side. This establishes the first inequality in 41, which completes the proof. ∎